OpenStudy (mayankdevnani):

Question below ::::::::::

OpenStudy (mayankdevnani):

A body of mass 6 kg is under a force of 6 N which causes displacement in it given by \[S=\frac{t^2}{4}m\].The work done by the force in 2 s is-

OpenStudy (mayankdevnani):

@amistre64 @ParthKohli @ajprincess @aroub @UnkleRhaukus @experimentX @ganeshie8 @gerryliyana @CliffSedge @mahmit2012

OpenStudy (anonymous):

Work = force × distance. Plug in your t=2s to solve for s, then multiply.

OpenStudy (mayankdevnani):

w=6*s

OpenStudy (mayankdevnani):

@phi have a look

OpenStudy (mayankdevnani):

\[S=\frac{2^2}{4}=1\]

OpenStudy (mayankdevnani):

right @ajprincess and @phi

OpenStudy (ajprincess):

yup perfect:)

OpenStudy (phi):

yes. s is 1 meter force is 6 N work is 6 N-m

OpenStudy (mayankdevnani):

sure @phi

OpenStudy (phi):

of course N-m is a joule so 6 joules

OpenStudy (mayankdevnani):

@phi

OpenStudy (phi):

Q14 does not say the force is 6 N.

OpenStudy (mayankdevnani):

it says 3J but you say 6J

OpenStudy (phi):

the question says "6 kg under a force that causes a displacement..." we need to find F, using F= ma we need a use d = 1/2 a t^2 we know d and t so we find a, then F and finally W = F d

OpenStudy (mayankdevnani):

@Aperogalics

OpenStudy (phi):

d (or s) in this case is 1 m t is 2 sec 1 = 1/2 a 2^2 a= 1/2 m/s^2 F= ma = 6*1/2 = 3 N W = F d = 3N * 1m = 3 N-m = 3 J

OpenStudy (mayankdevnani):

so my question is also same @phi

OpenStudy (phi):

your question says: mass 6 kg is under a force of 6 N ... you gave the force as 6 N so that is what I would use....

OpenStudy (mayankdevnani):

what do you mean? @phi

OpenStudy (phi):

Q14 requires that you find the force... Your 1st question may be inconsistent? i.e. 6 kg and 6 N don't mesh with the displacement?

OpenStudy (mayankdevnani):

oh!!

OpenStudy (mayankdevnani):

thnx..... @phi